### v1

@@ Discussion
Samples with high error values tend to be near the boundary of the training data space, where the model operates in an extrapolation regime. An example of this behavior in the HDD predictions is presented in Figure 9, where one test data sample is indicated by the crosshairs (dashed lines), which is zoomed in in Figure 10. The predictions for this particular sample are also shown in this plot. This sample demonstrates what we call model conservatism, which is an instance of a model’s predictions being biased and shifted towards the training data in case the test data sample is the extrapolation regime. This figure represents the predictions being shifted toward the test samples, which shows that the model is consistent with the data on which it was trained.
The same pattern of behavior is demonstrated in Figure 11. Here, the crosshairs are pointed on the test sample which is focused on in Figure 12 together with the predictions. It is observed that the predictions are located close to the test samples with the latter being in the extrapolation regime. This is another example of the model conservatism.
Conclusion
@@
  • Both parametric and nonparametric models demonstrate similar prediction qualities for mean IOPS and latency values. But the parametric model estimates the standard deviations better.
@@
Acknowledgments
The publication was supported by the grant for research centers in the field of AI provided by the Analytical Center for the Government of the Russian Federation (ACRF) in accordance with the agreement on the provision of subsidies (identifier of the agreement  [^00] D730321P5Q0002) and the agreement with HSE University No. 70-2021-00139.
The computation for this research was performed using the computational resources of HPC facilities at HSE University [26].

### v2

@@ VII Discussion
The results demonstrate distinct patterns in model performance based on input parameters. For instance, as shown in Figures 4 and 5, both models capture trends in IOPS and latency under varying queue depths and read/write fractions, with CatBoost showing consistently lower errors for standard deviation predictions. Additionally, the Normalizing Flow model’s conservative behavior in extrapolation scenarios, such as those depicted in Figure 9, highlights its tendency to bias predictions toward the training data in rare configurations. By combining insights from these patterns, we can better understand the trade-offs between precision and distribution modeling across configurations.
Samples with high error values tend to be near the boundary of the training data space, where the model operates in an extrapolation regime. An example of this behavior in the HDD predictions is presented in Figure 9(a), where one test data sample is indicated by the crosshairs (dashed lines), which is zoomed in in Figure 9(b). The predictions for this particular sample are also shown in this plot. This sample demonstrates what we call model conservatism, which is an instance of a model’s predictions being biased and shifted towards the training data in case the test data sample is the extrapolation regime. This figure represents the predictions being shifted toward the test samples, which shows that the model is consistent with the data on which it was trained.
The same pattern of behavior is demonstrated in Figure 9(c). Here, the crosshairs are pointed on the test sample which is focused on in Figure 9(d) together with the predictions. It is observed that the predictions are located close to the test samples with the latter being in the extrapolation regime. This is another example of the model conservatism.
VIII Conclusion
Storage system simulation models are indispensable tools in the design, optimization, and management of data storage systems. These models enable a detailed analysis of performance metrics such as throughput, latency, and I/O operations per second (IOPS), providing insights into how storage systems perform under various conditions. By replicating real-world workloads, simulation models help identify bottlenecks and optimize resource allocation. They serve as crucial instruments for designing and fine-tuning storage system architectures, allowing engineers to evaluate the impact of different configurations and workload scenarios. This leads to optimized system parameters that achieve desired performance goals. Furthermore, simulation models offer a cost-effective alternative to risky and expensive live environment changes, enabling extensive testing and validation of new technologies and configurations in a controlled setting. Their predictive analysis capabilities are essential for capacity planning, ensuring that storage systems can be efficiently scaled to meet future demands. Moreover, they facilitate detailed cost-benefit analysis, helping organizations evaluate the financial implications of different storage technologies and configurations, and assess the economic viability of upgrades and expansions. In addition to performance benchmarking, simulation models play a key role in energy efficiency by optimizing the storage system to minimize energy requirements.
@@ VIII Conclusion
  • Both parametric and nonparametric models demonstrate similar prediction qualities for mean IOPS and latency values. However, the parametric model estimates the standard deviations better.
@@
There are multiple constraints associated with the methods presented in this work. The models need data for training, and the accuracy of the predictions relies heavily on the size of the sample. It is necessary to employ multiple system configurations and generate a diverse range of data loads. The data collection process requires access to a real system and can be time-consuming. In addition, the models show sensitivity to the coverage of the input parameter space. Unreliable predictions can occur when the training sample lacks measurements from specific regions of configurations and data load parameter values. We can not rely on predictions for the external measurement regions because of the absence of relevant information in the training data.
There are several ways of future work for further improvement of the presented approach. The first is dedicated to increasing the liability of the models by estimating the uncertainties of the predictions. The predictions of the machine learning models are not ideal and the error values can be different for different inputs. Estimating prediction errors for each given data load parameter will help us decide whether we can rely on the prediction or not. The second direction is related to the explainability of the models. Currently, the models make performance predictions and do not provide an understanding of how the inputs affect the result. Information about how the prediction changes with changing the read fraction, queue depth, number of jobs, or other parameters will give a better understanding of the behavior of the system and will help optimize its performance.